Paolo owns a pizza shop. The price of pizza recently increased from $3 to $5 a slice. Paolo responded by increasing the quantity of slices he supplied from 100 to 150 slices per day. By the midpoint method, Paolo's price elasticity of supply is:
Answer #1

Answer is 0.8

Explanation: Mid-point elasticity formula is: (Q2 - Q1)/[(Q2 + Q1)/2] divided by (P2 - P1)/[(P2 + P1)/2]. Here Q2 and P2 are new quantity and price, and Q1 and P1 are old quantity and price. So elasticity = (150 - 100)/[(150 + 100)/2] divided by (5 - 3)/[(5 + 3)/2]. So 50/125 divided by 2/4. So (50)(4)/(125)(2) = 200/250 = 0.8

Please upvote this answer.